# How to Bridge Structural and Temporal Heterogeneity in Link Prediction? A Contrastive Method

### v1

@@ Abstract.
Temporal Heterogeneous Networks play a crucial role in capturing the dynamics and heterogeneity inherent in various real-world complex systems, rendering them a noteworthy research avenue for link prediction. However, existing methods fail to capture the fine-grained differential distribution patterns and temporal dynamic characteristics, which we refer to as spatial heterogeneity and temporal heterogeneity. To overcome such limitations, we propose a novel Contrastive Learning-based Link Prediction model, CLP, which employs a multi-view hierarchical self-supervised architecture to encode spatial and temporal heterogeneity. Specifically, aiming at spatial heterogeneity, we develop a structural feature modeling layer to capture the fine-grained topological distribution patterns from node- and edge-level representations, respectively. Furthermore, aiming at temporal heterogeneity, we devise a temporal information modeling layer to perceive the evolutionary dependencies of dynamic graph topologies from time-level representations. Finally, we encode the structural and temporal distribution heterogeneity from a contrastive learning perspective, enabling a comprehensive self-supervised hierarchical relation modeling for the link prediction task. Extensive experiments conducted on four real-world dynamic heterogeneous network datasets verify that our CLP consistently outperforms the state-of-the-art models, demonstrating an average improvement of 10.10%, 13.44% in terms of AUC and AP, respectively.
@@
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
Contemporary information networks such as social networks (Wang et al., 2024a) and biological systems (Chen and Liu, 2022) are becoming increasingly complex. These networks often comprise multi-typed nodes and connections, undergo continuous temporal evolution, making the link prediction in such complex networks a long-standing challenge. Specifically, the link prediction task aims to predict the likelihood of future connections between arbitrary nodes (Wang et al., 2024b; Yao and Li, 2024; Wu et al., 2022), which captures the evolution of heterogeneous networks and stores the temporal details of the node embeddings, simulating intricate and expressive semantics for real-world systems, including Social Recommendations (Wu et al., 2023a), Traffic Management (Shao et al., 2022; Jin et al., 2022), Medical Health (Wang et al., 2021b) and Network Biology (Guzzi and Zitnik, 2022). Aiming at modeling the dynamics and complex relationships between entities, link prediction models are primarily designed to portray the topological relationship between heterogeneous snapshots and the evolving progress along chronological order, revealing the distribution patterns within complex Temporal Heterogeneous Networks (THNs).
@@
The primary challenges in link prediction tasks revolve around heterogeneous entity relationship modeling and dynamic snapshot variability modeling. Current link prediction methods in literature typically segment dynamic snapshot sequences chronologically and address the complex entity relationships existing in each snapshot. We term such challenges as spatial complexity and temporal complexity. (1) Spatial complexity (Wang et al., 2019; Zhang et al., 2023d) highlights the complex heterogeneous static relationships between multi-typed entities in complex networks, primarily modeling the diverse co-occurrence paradigm through heterogeneous network embedding approaches. Specifically, Meta-Path-based approaches (Qiu et al., 2018; Dong et al., 2017; Fu et al., 2017; Zhao et al., 2023b) construct meta-paths within individual snapshots to excavate the heterogeneous information. On the other hand, Attribute-based methods (Hu et al., 2020a; Li et al., 2020; Zhang et al., 2023e) focus on incorporating multiple rich attributes (Xu et al., 2021; Peng et al., 2023) and merging neighbor attributes (Hu et al., 2020a; Li et al., 2020) to enhance the node embedding process. (2) Temporal complexity (Pareja et al., 2020; Pham et al., 2021; Fan et al., 2022; Layne et al., 2023) mainly exploits the dynamic distribution changes in snapshot sequences. Temporal approaches focus on tracking the continuous evolution across chronological snapshots, primarily classified into sequential and graph methods. Sequential methods (Hao et al., 2020; Jiang et al., 2023) learn from time-ordered snapshot sequences, capturing evolutionary dependencies between different snapshots based on Recurrent Neural Networks (RNNs) (Chen et al., 2019) and attention mechanism (Zhang et al., 2023c). Graph methods (Lei et al., 2019; Yang et al., 2019; Xu et al., 2020; Yang et al., 2022) aggregate embeddings of dynamic nodes, encoding the appearing or disappearing network features continuously over time through Graph Neural Networks (GNNs) (Zhao et al., 2023a; Gao et al., 2024).
@@
We illustrate such fine-grained differential distributions at the node-, edge-, and time-level in Figure 1. From the node-level propagation, taking the static THN snapshot $\mathcal{G}^{1}$ as an example, different propagation and aggregation paradigms convey differential information. If we take the quadrangle ([Uncaptioned image]) as the ego node, from the perspective of connection types, the propagation priority sequence for [Uncaptioned image] reveals: [Uncaptioned image] $=$ [Uncaptioned image] $=$ [Uncaptioned image]. Meanwhile, from the perspective of single-type neighbor numbers, the priority manifests: [Uncaptioned image] $>$ [Uncaptioned image] $=$ [Uncaptioned image], reflecting significant distribution heterogeneity among different propagation patterns. Analogously, targeting edge-level propagation, different edge types form the holistic propagation paradigm, yet convey unique propagation information. However, the heterogeneity between different edge-level propagation lacks sufficient attention. Finally, from time-level propagation, significant variations occur due to diverse inspection degrees (e.g., number of nodes, variations of edges), highlighting the transformative information conveyed by different time-level propagation patterns. We term such differential distribution in THNs as ‘spatial heterogeneity’ and ‘temporal heterogeneity’, which are crucial factors for link prediction modeling but have been rarely addressed in related research.
As is illustrated in the above example, different entities in complex networks possess variously-grained distribution differentiation, existing among nodes, edges and varying along chronological order. However, conventional methods ignore the modeling of such heterogeneity differentiation, which introduces the first challenge in this work, i.e., CH1: How to capture the heterogeneity differentiation existing in link relation networks? To represent such distribution heterogeneity in link prediction, it is essential to characterize the fine-grained intrinsic topological distribution in the link graph. Accordingly, we resort to Self-supervised Learning (Chen et al., 2020) to investigate the inherent inter-relation in temporal heterogeneous graphs.
@@ 2.1. Problem Formalization
$\mathbf{u}_{a}^{G}$ The temporal representation of node $a$ via Gated Recurrent Units (GRU).
$\mathbf{u}_{a}^{L}$ The temporal representation of node $a$ via Long Short-Term Memory (LSTM).
@@ 3. Methodology
The primary objective of our CLP is to learn a deep representation of a THN from various dynamic node and edge types to express the spatial discrepancy and temporal nonuniformity. To this end, we design CLP with a hierarchical architecture to capture the distribution heterogeneity, including (1) Structural Feature Modeling Layer, (2) Temporal Information Modeling Layer, and (3) Output Layer.
(1) Structural Feature Modeling Layer: First, we propose a two-layer hierarchical Graph Attention Network (GAT) to represent diverse types of edges and nodes within the THN from both node- and edge-level perspectives. Additionally, we introduce a contrastive representation method to differentiate feature heterogeneity at the node and edge levels, enhancing our ability to capture structural heterogeneity.
(2) Temporal Information Modeling Layer: Then, we deploy LSTM and GRU models to independently analyze temporal snapshot pattern, capturing the long-term and short-term dependencies between snapshots, respectively. Additionally, we implement contrastive learning strategies to bridge differences between these two sequence learning paradigms, thus preserving the temporal heterogeneity.
@@
3.1. Structural Feature Modeling Layer
The Structural Feature Modeling Layer aims to capture the structural distribution patterns and eliminate the discrepancy between nodes and edges in each static snapshot $\mathcal{G}^{t}\in\mathcal{G}$. Specifically, we initially partition the static snapshots $\mathcal{G}=\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots,\mathcal{G}^{T}\}$ into type-specific sub-networks based on different edge types. Then, we devise a two-layer hierarchical GAT to represent diverse types of nodes and edges within the THN from the perspectives of both node- and edge-levels in the Node-level Feature Learning Module and Edge-level Feature Learning Module, respectively. Additionally, we devise the contrastive representation heterogeneity differentiation modeling in the node- and edge- feature learning module to model the structural heterogeneity.
@@ 3.1.1. Node-level Feature Learning Module
[…]
@@
(3) $\mathbf{u}_{a}^{rt}=\sigma\left(\sum_{b\in\mathcal{N}_{a}^{rt}}\alpha_{a,b}^{rt}\mathbf{W}^{rt}\mathbf{x}_{b}\right),$
@@
Additionally, we implement a node-level InfoNCE loss function (van den Oord et al., 2018). This function guarantees that representations $\mathbf{u}_{a}^{rt}$ and $\mathbf{h}_{a}^{rt}$, derived from GAT and GNN for the same node $a$, are similar. Conversely, representation $\mathbf{u}_{b}^{rt}$, derived from GAT for a different node $b$, remains distinct. Within this framework, the pair ($\mathbf{u}_{a}^{rt}$, $\mathbf{h}_{a}^{rt}$) constitutes a positive sample pair, while the pair ($\mathbf{u}_{a}^{rt}$, $\mathbf{u}_{b}^{rt}$) functions as a negative sample pair, as specified in Eq. (5) and Eq. (6):
(5) $\mathcal{L}^{{+}}_{{Node}}=-\sum_{t=1}^{T}\sum_{r=1}^{R}\sum_{a=1}^{\left|V^{t}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{h}_{a}^{rt}\right)/\tau\right)}{\sum_{b\in\mathcal{N}_{a}^{rt}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{h}_{b}^{rt}\right)/\tau\right)},$
(6) $\mathcal{L}^{{-}}_{{Node}}=-\sum_{t=1}^{T}\sum_{r=1}^{R}\sum_{a=1}^{\left|V^{t}\right|}\sum_{\begin{subarray}{c}b\in\mathcal{N}_{a}^{rt}\land b\neq a\end{subarray}}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{u}_{b}^{rt}\right)/\tau\right)}{\sum_{k\in\mathcal{N}_{a}^{rt}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{u}_{k}^{rt}\right)/\tau\right)},$
@@ 3.1.2. Edge-level Feature Learning Module
The Node-level Feature Learning module captures information specific to one edge type. However, heterogeneous networks typically comprise multiple edge types. To equip the information from all edge types at each node, we devise the Edge-level Feature Learning module, which determines the importance weights for different edge types. This module aggregates different forms of information for a specific type, thereby generating the node’s embedding enriched with heterogeneous edge information. Specifically, each node’s embedding vector undergoes a nonlinear mapping. The attention weight $\delta_{a}^{rt}$ between each edge type $r$ and node $a$ in the $t$-th snapshot graph is obtained through the softmax activation function, as described in Eq.(7) and Eq. (8):
@@
(8) $\delta_{a}^{rt}=\frac{\exp\left(\gamma_{a}^{rt}\right)}{\sum_{r^{\prime}\in R}\exp\left(\gamma_{a}^{r^{\prime}t}\right)},$
@@
(10) $\mathbf{h}_{a}^{t}=\frac{1}{R}\sum_{r=1}^{R}\left(\mathbf{u}_{a}^{rt}+\frac{1}{\sqrt{\left|\mathcal{N}_{a}^{rt}\right|\cdot\left|\mathcal{N}_{b}^{rt}\right|}}\sum_{b\in\mathcal{N}_{a}^{rt}}\mathbf{u}_{b}^{rt}\right),$
@@
Furthermore, to confirm that the edge-specific embedding $\mathbf{u}_{a}^{t}$ for node $a$ is similar to its embedding $\mathbf{h}_{a}^{t}$ in the unified graph and dissimilar to the aggregated representation $\mathbf{u}_{b}^{t}$ for node $b$, we define the edge-level InfoNCE loss function considering the heterogeneous and unified aggregated representations of the same node as a positive sample pair ($\mathbf{u}_{a}^{t}$, $\mathbf{h}_{a}^{t}$), while forming a negative sample pair ($\mathbf{u}_{a}^{t}$, $\mathbf{u}_{b}^{t}$) with heterogeneous aggregated representations from different nodes, as Eq. (11) and Eq. (12):
(11) $\mathcal{L}^{{+}}_{{Edge}}=-\sum_{t=1}^{T}\sum_{a=1}^{\left|V^{t}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{h}_{a}^{t}\right)/\tau\right)}{\sum_{b\in\mathcal{N}_{a}^{t}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{h}_{b}^{t}\right)/\tau\right)},$
(12) $\mathcal{L}^{{-}}_{{Edge}}=-\sum_{t=1}^{T}\sum_{a=1}^{\left|V^{t}\right|}\sum_{\begin{subarray}{c}b\in\mathcal{N}_{a}^{t}\land b\neq a\end{subarray}}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{u}_{b}^{t}\right)/\tau\right)}{\sum_{k\in\mathcal{N}_{a}^{t}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{u}_{k}^{t}\right)/\tau\right)},$

### v3

# How to Bridge Spatial and Temporal Heterogeneity in Link Prediction? A Contrastive Method

@@ Abstract.
Temporal Heterogeneous Networks play a crucial role in capturing the dynamics and heterogeneity inherent in various real-world complex systems, rendering them a noteworthy research avenue for link prediction. However, existing methods fail to capture the fine-grained differential distribution patterns and temporal dynamic characteristics, which we refer to as spatial heterogeneity and temporal heterogeneity. To overcome such limitations, we propose a novel Contrastive Learning-based Link Prediction model, CLP, which employs a multi-view hierarchical self-supervised architecture to encode spatial and temporal heterogeneity. Specifically, aiming at spatial heterogeneity, we develop a spatial feature modeling layer to capture the fine-grained topological distribution patterns from node- and edge-level representations, respectively. Furthermore, aiming at temporal heterogeneity, we devise a temporal information modeling layer to perceive the evolutionary dependencies of dynamic graph topologies from time-level representations. Finally, we encode the spatial and temporal distribution heterogeneity from a contrastive learning perspective, enabling a comprehensive self-supervised hierarchical relation modeling for the link prediction task. Extensive experiments conducted on four real-world dynamic heterogeneous network datasets verify that our CLP consistently outperforms the state-of-the-art models, demonstrating an average improvement of 10.10%, 13.44% in terms of AUC and AP, respectively.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Networks Network dynamics
@@ 1. Introduction
Contemporary information networks such as social networks (WangLH24) and biological systems (chen2022graph) are becoming increasingly complex. These networks often comprise multi-typed nodes and connections, undergo continuous temporal evolution, making the link prediction in such complex networks a long-standing challenge. Specifically, the link prediction task aims to predict the likelihood of future connections between arbitrary nodes (WangLSCFLG24; YaoL24; WuSGG22), which captures the evolution of heterogeneous networks and stores the temporal details of the node embeddings, simulating intricate and expressive semantics for real-world systems, including Social Recommendations (wu2023pda), Traffic Management (JinRZJGLMW22), Medical Health (wang2021multi) and Network Biology (GuzziZ22). Aiming at modeling the dynamics and complex relationships between entities, link prediction models are primarily designed to portray the topological relationship between heterogeneous snapshots and the evolving progress along chronological order, revealing the distribution patterns within complex Temporal Heterogeneous Networks (THNs).
@@
The primary challenges in link prediction tasks revolve around heterogeneous entity relationship modeling and dynamic snapshot variability modeling. Current link prediction methods in literature typically segment dynamic snapshot sequences chronologically and address the complex entity relationships existing in each snapshot. We term such challenges as spatial complexity and temporal complexity. (1) Spatial complexity (wang2019heterogeneous; zhang2023page) highlights the complex heterogeneous static relationships between multi-typed entities in complex networks, primarily modeling the diverse co-occurrence paradigm through heterogeneous network embedding approaches. Specifically, Meta-Path-based approaches (qiu2018network; dong2017metapath2vec; fu2017hin2vec; zhao2023link) construct meta-paths within individual snapshots to excavate the heterogeneous information. On the other hand, Attribute-based methods (hu2020heterogeneous; li2020type; zhang2023dynamic1) focus on incorporating multiple rich attributes (xu2021topic; peng2023th) and merging neighbor attributes (hu2020heterogeneous; li2020type) to enhance the node embedding process. (2) Temporal complexity (pareja2020evolvegcn; pham2021comgcn; fan2022heterogeneous) mainly exploits the dynamic distribution changes in snapshot sequences. Temporal approaches focus on tracking the continuous evolution across chronological snapshots, primarily classified into sequential and graph methods. Sequential methods (hao2020dynamic; jiang2023dnformer) learn from time-ordered snapshot sequences, capturing evolutionary dependencies between different snapshots based on Recurrent Neural Networks (RNNs) (chen2019lstm) and attention mechanism (zhang2023attentional). Graph methods (lei2019gcn; yang2019advanced; XuRKKA20; yang2022few) aggregate embeddings of dynamic nodes, encoding the appearing or disappearing network features continuously over time through Graph Neural Networks (GNNs).
@@
We illustrate such fine-grained differential distributions at the node-, edge-, and time-level in Figure 1. From the node-level propagation, taking the static THN snapshot $\mathcal{G}^{1}$ as an example, different propagation and aggregation paradigms convey differential information. If we take the quadrangle ([Uncaptioned image]) as the ego node, from the perspective of connection types, the propagation priority sequence for [Uncaptioned image] reveals: [Uncaptioned image] == [Uncaptioned image] == [Uncaptioned image]. Meanwhile, from the perspective of single-type neighbor numbers, the priority manifests: [Uncaptioned image] >> [Uncaptioned image] == [Uncaptioned image], reflecting significant distribution heterogeneity among different propagation patterns. Analogously, targeting edge-level propagation, different edge types form the holistic propagation paradigm, yet convey unique propagation information. However, the heterogeneity between different edge-level propagation lacks sufficient attention. Finally, from time-level propagation, significant variations occur due to diverse inspection degrees (e.g., number of nodes, variations of edges), highlighting the transformative information conveyed by different time-level propagation patterns. We term such differential distribution in THNs as ‘spatial heterogeneity’ and ‘temporal heterogeneity’, which are crucial factors for link prediction modeling but have been rarely addressed in related research.
As is illustrated in the above example, different entities in complex networks possess variously-grained distribution differentiation, existing among nodes, edges and varying along chronological order. However, conventional methods ignore the modeling of such heterogeneity differentiation, which introduces the first challenge in this work, i.e., CH1: How to capture the heterogeneity differentiation existing in link relation networks? To represent such distribution heterogeneity in link prediction, it is essential to characterize the fine-grained intrinsic topological distribution in the link graph. Accordingly, we resort to Self-supervised Learning (chen2020simple) to investigate the inherent inter-relation in temporal heterogeneous graphs.
@@ 2.1. Problem Formalization
$\mathbf{u}_{a}^{L}$ The long-term temporal representation of node $a$.
$\mathbf{u}_{a}^{S}$ The short-term temporal representation of node $a$.
@@ 3. Methodology
The primary objective of our CLP is to learn a deep representation of a THN from various dynamic node and edge types to express the spatial discrepancy and temporal nonuniformity. To this end, we design CLP with a hierarchical architecture to capture the distribution heterogeneity, including (1) Spatial Feature Modeling Layer, (2) Temporal Information Modeling Layer, and (3) Output Layer.
(1) Spatial Feature Modeling Layer: First, we propose a two-layer hierarchical Graph Attention Network (GAT) to represent diverse types of edges and nodes within the THN from both node- and edge-level perspectives. Additionally, we introduce a contrastive representation method to differentiate feature heterogeneity at the node and edge levels, enhancing our ability to capture spatial heterogeneity.
(2) Temporal Information Modeling Layer: Then, we deploy Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU) models to independently analyze temporal snapshot pattern, capturing the long-term and short-term dependencies between snapshots, respectively. Additionally, we implement contrastive learning strategies to bridge differences between these two sequence learning paradigms, thus preserving the temporal heterogeneity.
@@
3.1. Spatial Feature Modeling Layer
The Spatial Feature Modeling Layer aims to capture the structural distribution patterns and eliminate the discrepancy between nodes and edges in each static snapshot $\mathcal{G}^{t}\in\mathcal{G}$. Specifically, we initially partition the static snapshots $\mathcal{G}=\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots,\mathcal{G}^{T}\}$ into type-specific sub-networks based on different edge types. Then, we devise a two-layer hierarchical GAT to represent diverse types of nodes and edges within the THN from the perspectives of both node- and edge-levels in the Node-level Feature Learning Module and Edge-level Feature Learning Module, respectively. Additionally, we devise the contrastive representation heterogeneity differentiation modeling in the node- and edge- feature learning module to model the spatial heterogeneity.
@@ 3.1.1. Node-level Feature Learning Module
where $\mathbf{x}_{a}$ and $\mathbf{x}_{b}$ are used to initialize node $a$ and $b$; $\mathbf{A}^{rt}$ and $\mathbf{W}^{rt}$ represent the learnable attention weight vector and mapping matrix, specific to $r$ type subgraphs of $\mathcal{G}^{t}$; The function $\sigma\left(\cdot\right)$ denotes the activation function and $\|$ signifies the concatenation operation. The formalized attention weight parameter $\alpha^{rt}_{ij}$ between node $a$ and $b$ in the $r$ type subgraph of $\mathcal{G}^{t}$ is defined as follows:
@@
(3) $\mathbf{u}_{a}^{rt}=\sigma\left(\sum_{b\in\mathcal{N}_{a}^{rt}}\alpha_{a,b}^{rt}\mathbf{W}^{rt}\mathbf{x}_{b}\right).$
@@
Additionally, we implement a node-level InfoNCE loss function (abs-1807-03748). This function guarantees that representations $\mathbf{u}_{a}^{rt}$ and $\mathbf{h}_{a}^{rt}$, derived from GAT and GNN for the same node $a$, are similar. Conversely, representation $\mathbf{u}_{b}^{rt}$, derived from GAT for a different node $b$, remains distinct. Within this framework, the pair ($\mathbf{u}_{a}^{rt}$, $\mathbf{h}_{a}^{rt}$) constitutes a positive sample pair, while the pair ($\mathbf{u}_{a}^{rt}$, $\mathbf{u}_{b}^{rt}$) functions as a negative sample pair, as specified in Eq. (5):
$\begin{aligned} \mathcal{L}^{{+}}_{{Node}}=-\sum_{t=1}^{T}\sum_{r=1}^{R}\sum_{a=1}^{\left|V^{t}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{h}_{a}^{rt}\right)/\tau\right)}{\sum_{b\in\mathcal{N}_{a}^{rt}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{h}_{b}^{rt}\right)/\tau\right)},\\ \mathcal{L}^{{-}}_{{Node}}=-\sum_{t=1}^{T}\sum_{r=1}^{R}\sum_{a=1}^{\left|V^{t}\right|}\sum_{\begin{subarray}{c}b\in\mathcal{N}_{a}^{rt}\land b\neq a\end{subarray}}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{u}_{b}^{rt}\right)/\tau\right)}{\sum_{k\in\mathcal{N}_{a}^{rt}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{rt},\mathbf{u}_{k}^{rt}\right)/\tau\right)},\end{aligned}$
@@ 3.1.2. Edge-level Feature Learning Module
The Node-level Feature Learning module captures information specific to one edge type. However, heterogeneous networks typically comprise multiple edge types. To equip the information from all edge types at each node, we devise the Edge-level Feature Learning module, which determines the importance weights for different edge types. This module aggregates different forms of information for a specific type, thereby generating the node’s embedding enriched with heterogeneous edge information. Specifically, each node’s embedding vector undergoes a nonlinear mapping. The attention weight $\delta_{a}^{rt}$ between each edge type $r$ and node $a$ in the $t$-th snapshot graph is obtained through the softmax activation function, as described in Eq.(6):
@@
$\delta_{a}^{rt}$ $={\exp\left(\gamma_{a}^{rt}\right)}/{\sum\nolimits_{r^{\prime}\in R}\exp\left(\gamma_{a}^{r^{\prime}t}\right)},$
@@
(8) $\mathbf{h}_{a}^{t}=\frac{1}{R}\sum_{r=1}^{R}(\mathbf{u}_{a}^{rt}+\frac{1}{\sqrt{|\mathcal{N}_{a}^{rt}|\cdot|\mathcal{N}_{b}^{rt}|}}\sum_{b\in\mathcal{N}_{a}^{rt}}\mathbf{u}_{b}^{rt}),$
@@
Furthermore, to confirm that the edge-specific embedding $\mathbf{u}_{a}^{t}$ for node $a$ is similar to its embedding $\mathbf{h}_{a}^{t}$ in the unified graph and dissimilar to the aggregated representation $\mathbf{u}_{b}^{t}$ for node $b$, we define the edge-level InfoNCE loss function considering the heterogeneous and unified aggregated representations of the same node as a positive sample pair ($\mathbf{u}_{a}^{t}$, $\mathbf{h}_{a}^{t}$), while forming a negative sample pair ($\mathbf{u}_{a}^{t}$, $\mathbf{u}_{b}^{t}$) with heterogeneous aggregated representations from different nodes, as expressed in Eq. (9):
$\begin{aligned} \mathcal{L}^{{+}}_{{Edge}}&=-\sum_{t=1}^{T}\sum_{a=1}^{\left|V^{t}\right|}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{h}_{a}^{t}\right)/\tau\right)}{\sum_{b\in\mathcal{N}_{a}^{t}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{h}_{b}^{t}\right)/\tau\right)},\\ \mathcal{L}^{{-}}_{{Edge}}&=-\sum_{t=1}^{T}\sum_{a=1}^{\left|V^{t}\right|}\sum_{\begin{subarray}{c}b\in\mathcal{N}_{a}^{t}\land b\neq a\end{subarray}}\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{u}_{b}^{t}\right)/\tau\right)}{\sum_{k\in\mathcal{N}_{a}^{t}}\exp\left(\operatorname{sim}\left(\mathbf{u}_{a}^{t},\mathbf{u}_{k}^{t}\right)/\tau\right)},\end{aligned}$